OpenStudy (anonymous):

log(2x^2+3x)=log(10x+30)

OpenStudy (raden):

first, take the equation : 2x^2 + 3x = 10x+30 2x^2 + 3x - 10x - 30 = 0 2x^2 - 7x - 30 = 0 solve for x

OpenStudy (mathstudent55):

If \( \log a = \log b\), then \(a = b\). That is what allows you to set \(2x^2 + 3x\) equal to \(10x + 30\). Now you have a quadratic in x. Move all terms to the left side and set equal to zero and solve the quadratic.